OpenStudy (anonymous):

If a line with a slope of -2 crosses the y-axis at (0, 3), what is the equation of the line?

OpenStudy (anonymous):

@Nnesha do you know?

Nnesha (nnesha):

okay so here is a equation y = mx + b where m is slope and b is y-intercept so replace m by -2

Nnesha (nnesha):

your y-intercept at (0,3) where x is zero so y-intercept is 3 point on y-axis is y-intercept

OpenStudy (calculusfunctions):

You already have the slope and the y-intercept. Thus the equation of a line in slope intercept form is y = (slope)x + y-intercept

OpenStudy (anonymous):

y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3 Those are my options there not in the same form

Nnesha (nnesha):

option are in standard form so what is your equation should be with -2 slope and 3 y-intercept

OpenStudy (anonymous):

y=-2x+3?

Nnesha (nnesha):

yes now add 2x both side

OpenStudy (anonymous):

would it be y+2x=3?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

David is rowing a boat upstream. The river is flowing at a speed of 2 miles per hour. David starts rowing at a speed of 6 miles per hour, and his speed decreases at a rate of 1.5 miles per hour. which equation gives the speed of the boat after David has rowed for x hours A.Y=4+1.5X B.Y=4-1.5X C.6-1.5X